OpenStudy (butterflyhope):

Solve: -9z - 36 - 6 = 48 I need alot of help today xD Thanks! :)

OpenStudy (anonymous):

\[-9z-42=48\]then \[-9z=90\] so, \[z=-10\]

OpenStudy (butterflyhope):

Can you work it out fully please? Thank you :)

OpenStudy (anonymous):

first: simplify the expression by adding the known together and the unknown together in each side\[-9z+(-36-6))=48\].\[-9z-42=48\] second: put the known together and the unknown together\[-9z=90\] third: you divide by the multiplier of the unknown both sides \[\frac{-9z}{-9}=\frac{90}{-9}\] the result is \[z=-10\]